What number whose product with 3 is the same as it's sum with 30?
Answer by cesardelrosario(3)

solution:
Let x=the number
y=the product
z=the sum
(x)(3) = y equation 1
x + 30 = z equation 2
but y = z, therefore 3x = x + 30
3x - x = 30
2x = 30
x = 30/2
x = 15